OpenStudy (anonymous):

5 questions, can some please check them

OpenStudy (anonymous):

1: what is the equation of the line? A: y = -3/2x B: y = -2/3x C: y = 2/3x D: y = 3/2x I got C

OpenStudy (anonymous):

2: Points A and B lie on the graph of the equation y = −4.3x. What are the y-coordinates for each point? I got -3, 12.9 and 5, 21.5 3:Choose the correct answer. Determine whether the given has a positive slope, a negative slope, a zero slope, or an undefined slope. A. negative slope B. undefined slope C. zero slope D. I got zero slope positive slope

OpenStudy (anonymous):

4: What is the slope of the line passing through the points (1, –5) and (4, 1)? i got 2

OpenStudy (anonymous):

5:Which is the graph of y=1/2x

OpenStudy (anonymous):

I got BB for the 5th one

OpenStudy (mathmate):

1. C is correct 2. Wrong sign for one of the values, values are correct 3. Zero slope is correct 4. 2 is correct 5. Nope, wrong sign again!

OpenStudy (anonymous):

so, what did I do wrong on 5 and 2

OpenStudy (mathmate):

I am sure you'll find the problems if you redo the calculations.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I think 5 is Cc

OpenStudy (mathmate):

cc is correct for 5, since it is a positive slope.

OpenStudy (mathmate):

Remember, when the slope is negative, such as y=-2x, x and y have different signs, unless x=y=0.

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

I wasn't sure for number 2 to put decimal points or commas

OpenStudy (mathmate):

If you are in the US, points are usually expressed as (-3, 12.9) for a point. In Europe, decimals are replaced by commas.

OpenStudy (anonymous):

I'm in us

OpenStudy (mathmate):

So the form (-3, 12.9) would be correct for the first point. You just have to make corrections for the second point.

OpenStudy (anonymous):

Okay

OpenStudy (mathmate):

Did you find the correction needed?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ty very much :)